The number of calories in a piece of pie is 20 less than three times the number of calories in a scoop of ice cream. The pie and ice cream together have 500 calories. How many calories are in each?
Answer by Cromlix(4381)

Hi there,
Make number of calories in a
scoop of ice cream = 'x' calories
Calories in a scoop of pie = 3x - 20
Total of scoop of pie + ice cream = 500 cals
x + (3x - 20) = 500
x + 3x - 20 = 500
Collect like terms
4x = 500 + 20
4x = 520
x = 130
Ice Cream has 130 Calories
Scoop of pie has 370 Calories.
Hope this helps :-)